40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Sep 29 15:58:19 2019, gautam, Update, LSC, POX locking attempt POXlockAttempt.png
    Reply  Mon Sep 30 14:16:43 2019, gautam, Update, LSC, POX PD checkout - solved 
Message ID: 14910     Entry time: Sun Sep 29 15:58:19 2019     Reply to this: 14915
Author: gautam 
Type: Update 
Category: LSC 
Subject: POX locking attempt 

Summary:

There is no visible PDH error signal on the POX11 channels. As a result, I am unable to lock the XARM length to the laser frequency. See Attachment #1 - the Y arm length is locked to the PSL frequency, and control is disabled for the XARM servo.

Details:

Now that several of the c1iscaux functionality tests have been completed, I wanted to push ahead with some locking. However, I was foiled at this early stage, for reasons as yet unknown. One possibility is that the

  • I am able to see TRX cavity flashes >0.8, which suggest to me that the cavity is well aligned.
  • Moreover, I am able to lock some (admittedly high TEM order) mode of the green laser, which further supports the above hypothesis.
  • However, there are no visible PDH-like features in the POX11_I or POX11_Q channels.
  • I checked that the cables from the output of the POX11 demod board are in fact going to the correct channels on WF1 (#5 and #6 respectively), and that the whitening gain for this channel is set to the nominal +30 dB.
  • Next, I went to the POX table and looked for the POX IR beam. I couldn't see anything, but this beam is expected to be weak (of the order of 1 W * T_PRM * R_AR_ITM ~ 30 uW), which is probably not so easily visible.

Next steps:

  • Look for the POX beam with an IR viewer.
  • Confirm that everything is order on the LSC Demod board for POX 11 - maybe the LO isn't connected (somehow)?
Attachment 1: POXlockAttempt.png  122 kB  Uploaded Sun Sep 29 17:06:58 2019  | Hide | Hide all
POXlockAttempt.png
ELOG V3.1.3-